In a straight line motion the distance travelled is propertional to the square root of the time taken. The acceleration of the particle is proportional to:

(a) velocity            (b) V2    

(c) V3                   (d) √V

How do we answer this question?

Grade:12

1 Answers

Paarth Arkadi
8 years ago
\alpha t½
By diffrentiation ,
\alpha 1/t1/2
Again by diffrentiation,
\alpha 1/ t3/2
Therefore a \alpha v3